# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio

### v2

@@ Abstract
Solving large-scale, continuous-time portfolio optimization problems involving numerous assets and state-dependent dynamics has long been challenged by the curse of dimensionality. Traditional dynamic programming and PDE-based methods, while rigorous, typically become computationally intractable beyond few state variables ($\sim$3-6 limit in prior studies). To overcome this critical barrier, we introduce the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework. PG-DPO leverages Pontryagin’s Maximum Principle (PMP) and backpropagation-through-time (BPTT) to guide neural network policies, handling exogenous states without dense grids. This PMP-guided approach holds potential for a broad class of sufficiently regular continuous-time control problems. Crucially, our computationally efficient “Two-Stage” variant exploits rapidly stabilizing BPTT costate estimates, converting them into near-optimal Pontryagin controls after only a short warm-up, significantly reducing training overhead. This enables a breakthrough in scalability: numerical experiments show PG-DPO successfully tackles problems with dimensions previously considered far out of reach (up to 50 assets and 10 state variables). The framework delivers near-optimal policies, offering a practical and powerful alternative for high-dimensional continuous-time portfolio choice.
@@ 1 Introduction
High-dimensional optimal control problems often fall prey to the “curse of dimensionality” (Bellman, , 1966), whereby the computational cost of classical dynamic programming (DP) approaches grows exponentially with the number of state variables. In a continuous-time finance setting, this difficulty becomes starkly evident in multi-asset Merton problems (Merton, , 1971), where an investor dynamically allocates wealth among numerous risky assets whose drift and volatility may depend on exogenous state variables such as macroeconomic indicators or stochastic volatility factors. While DP or Hamilton–Jacobi–Bellman (HJB) PDE solvers provide a rigorous framework for low-dimensional cases (Bertsekas & Tsitsiklis, , 1995), they become intractable as the number of assets or state dimensions increases, with prior influential studies often limited to analyzing problems involving up to approximately 3-5 state variables (Campbell & Viceira, , 1999; Lynch, , 2001; Balduzzi & Lynch, , 1999) or specific frameworks handling around 6 state variables (Garlappi & Skoulakis, , 2010), often necessitating simulation-based approaches for tractability (Brandt et al., , 2005).
@@
A promising alternative is to employ Pontryagin’s Maximum Principle (PMP), an indirect method that bypasses the explicit computation of value functions (Pontryagin, , 2014; Yong & Zhou, , 2012) Rather than storing or approximating a high-dimensional value grid, the PMP characterizes the optimal control in terms of forward–backward differential equations for the state and its costate (adjoint). Notably, adding one more state dimension does not produce an exponential explosion in computational requirements, in contrast to DP (Bellman, , 1966; Campbell et al., , 2003). This makes Pontryagin-based methods appealing for large-scale portfolio optimization problems involving many assets and exogenous state processes (Karatzas et al., , 1987).
@@
Recent advances in deep learning further enhance the scalability of methods for solving high-dimensional stochastic control problems (e.g., Han et al., , 2018; Raissi et al., , 2019, for general PDE-based approaches). Concurrently, deep learning techniques are being applied directly to financial dynamic programming. Some approaches tackle infinite horizon problems via the HJB equation (Duarte et al., , 2024), while others aim to directly solve the associated nonlinear PDEs using methods like deep branching algorithms (Nguwi et al., , 2024). However, applications such as the Merton HJB have revealed limitations in accuracy or efficiency for certain PDE solvers like Deep BSDE variants (Nguwi et al., , 2024). Our PMP-based PG-DPO framework offers a distinct path, naturally handling finite horizons and focusing on the optimality conditions derived from PMP, which are applicable to a wide range of stochastic control problems possessing sufficient regularity. One notable strategy related to the use of adjoint methods in optimal control is to directly parameterize the policy (i.e., consumption and investment controls) with neural networks, then apply backpropagation-through-time (BPTT) to enforce Pontryagin’s optimality conditions (Huh, , 2024). This strategy, known as “Pontryagin-Guided Direct Policy Optimization (PG-DPO),” circumvents the need to approximate value functions or PDE solutions. Instead, it uses simulation-based gradients that automatically incorporate the adjoint processes, drastically reducing computational overhead in high-dimensional settings (Kushner & Yin, , 2003; Borkar & Borkar, , 2008).
@@
Furthermore, we introduce a “Two-Stage” PG-DPO variant that exploits the observation that costate estimates often stabilize much faster than the full policy network. Once the costates are sufficiently stable (after a short warm-up), one can bypass the potentially suboptimal neural policy outputs and directly substitute the costate estimates into the PMP first-order conditions. This recovers a near-optimal control analytically at every time-state point with minimal additional training, significantly reducing computational burden while maintaining high accuracy. This efficiency enables the solution of problems with dozens of assets and state variables (e.g., n=50, k=10 in our experiments), a scale considerably larger than those typically addressed by the numerical methods cited previously. This demonstrated scalability on a benchmarkable problem highlights the potential of leveraging PMP structure via BPTT estimates for tackling other complex, high-dimensional problems that may exhibit similar amenability to costate estimation under suitable regularity conditions.
@@
  • We introduce Pontryagin-Guided Direct Policy Optimization (PG-DPO), featuring a highly efficient “Two-Stage” variant that uses PMP structure and BPTT-derived costates to solve complex continuous-time control problems.
  • We demonstrate through numerical experiments the framework’s ability to break the dimensional barrier, successfully tackling portfolio optimization problems with dozens of assets and state factors (up to n=50, k=10).
In the remainder of this paper, we first present the multi-asset Merton problem with exogenous states in Section 2. Section 3 introduces our Pontryagin-Guided Direct Policy Optimization (PG-DPO) algorithm and its Two-Stage extension. Numerical experiments presented in Section 4 confirm the framework’s scalability and near-optimal performance within this testbed, illustrating its potential for high-dimensional applications. Finally, Section 5 concludes and discusses potential future directions.
@@ 2 Multi-Asset Continuous-Time Portfolio Problem with State Variables
In this section, we present a multi-asset extension of Merton’s continuous-time portfolio problem (Merton, , 1971; Kim & Omberg, , 1996; Liu, , 2007), allowing the investor to allocate wealth across multiple (potentially correlated) risky assets and a risk-free asset, while consuming continuously. Each risky asset features its own drift and volatility, leading to an $n$-dimensional portfolio weight vector $\bm{\pi}_{t}$. Furthermore, an exogenous state process $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ influences market parameters such as returns and interest rates, adding additional complexity to the investor’s decision.
Although the wealth $X_{t}$ remains scalar, the investor’s optimal policy now depends on both $\bm{\pi}_{t}\in\mathbb{R}^{n}$ and a high-dimensional state $\mathbf{Y}_{t}$. We handle this through Pontryagin’s maximum principle (Pontryagin, , 2014; Yong & Zhou, , 2012), introducing a costate for $X_{t}$ and another for $\mathbf{Y}_{t}$. Numerically, one can represent $\{\bm{\pi}_{t},C_{t}\}$ via neural networks, thereby enabling policy gradient or dynamic programming algorithms even in multi-asset, state-dependent environments.
@@ 2.1 Problem Formulation with State Variables and Neural Network Controls
where $\bm{\mu}_{Y}(t,\mathbf{Y}_{t})$ is the $k\times 1$ drift vector, $\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})$ is often assumed to be a $k\times k$ diagonal matrix of diffusion coefficients, and $\mathbf{W}^{Y}_{t}$ is a $k\times 1$ vector of potentially correlated Brownian motions driving the state variables.
@@
    Boundedness and proportional consumption. There exist constants $\Pi_{\max}>0$ and $\xi\in(0,1)$ such that
@@
    Positive wealth. The corresponding wealth process $X_{t}$ solving (1) remains strictly positive almost surely:
@@
The investor’s objective is to choose an admissible control pair $(\bm{\pi},C)$ to maximize the expected discounted utility functional $J$:
[…]
subject to the wealth dynamics (1) and admissibility constraints (Assumption 2). Here $\delta>0$ is the subjective discount rate, and $\kappa\geq 0$ weights the utility of terminal wealth (bequest motive). The utility function $U:(0,\infty)\to\mathbb{R}$ represents the investor’s preferences. A widely used specification is the Constant Relative Risk Aversion (CRRA) form:
$U(x)=\begin{cases}\displaystyle\frac{x^{1-\gamma}}{1-\gamma},&\gamma>0,\gamma\neq 1,\\[6.0pt] \ln(x),&\gamma=1,\end{cases}$
with $\gamma>0$ denoting the coefficient of relative risk aversion (see Pennacchi, , 2008).
Assumption 2 (iii) bounds consumption by $C_{t}\leq\xi X_{t}$ with $\xi\in(0,1)$, while Assumption 4 (b) gives a $C^{2}$, strictly concave utility $U$ satisfying the Inada limits. Because $\delta>0$ and $X_{t}>0$ a.s. (Assumption 2 (iv)), we have
$\mathbb{E}\!\bigl{[}\,e^{-\delta t}\,U(C_{t})\bigr{]}\leq e^{-\delta t}\,U\!\bigl{(}\xi\,\mathbb{E}[X_{t}]\bigr{)}<\infty,\quad\forall t\in[0,T],$
and similarly $\mathbb{E}[U(X_{T})]<\infty$. Hence the functional $J(\bm{\pi},C)$ in (2) is finite for every admissible control pair, so the value function $V$ is well defined.
@@
Additional assumptions are needed for the well-posedness of the problem and regularity of the value function.
@@
    Uniform non-degeneracy of risky-asset volatility. Let $a(t,\mathbf{y}):=\bm{\sigma}(t,\mathbf{y})\,\Psi\,\bm{\sigma}(t,\mathbf{y})^{\top}$ be the instantaneous covariance matrix of the risky asset returns (Note: original used $\bm{\sigma}\bm{\sigma}^{\top}$, assuming $\Psi=I$; using $\bm{\sigma}\Psi\bm{\sigma}^{\top}$ is more general if $\mathbf{W}^{X}$ components are correlated and $\bm{\sigma}$ is diagonal). We assume there exists $\lambda>0$ such that
@@
    (Equivalently, the covariance matrix of risky returns $a(t,\mathbf{y})$ is uniformly positive definite, ensuring invertibility, which is crucial for solving the portfolio problem.)
@@
    Utility specification. The utility function $U:(0,\infty)\to\mathbb{R}$ satisfies $U\in C^{2}(0,\infty)$, $U^{\prime}>0,\;U^{\prime\prime}<0$ (strictly increasing and concave), and the Inada conditions $\lim_{c\downarrow 0}U^{\prime}(c)=\infty,\;\lim_{c\uparrow\infty}U^{\prime}(c)=0.$
@@
    Discount and bequest parameters. The subjective discount rate satisfies $\delta>0$ and the terminal weight satisfies $\kappa\geq 0$.
Assumption 4 (a) guarantees the diffusion related to risky assets is non-degenerate, a key ingredient for the regularity of the associated Hamilton-Jacobi-Bellman (HJB) PDE. Parts (b)-(c) ensure the objective functional is well-defined and satisfies standard preference assumptions. Combined with the coefficient regularity in Assumption 1, these conditions typically yield the existence, uniqueness, and sufficient smoothness (e.g., $V\in C^{1,2,2}$) of the solution to the HJB equation, allowing for verification theorems to be applied; see Oksendal, (2013, Chs. 6-7) and related literature on optimal portfolio choice.

### v3

@@ Abstract
Solving large-scale, continuous-time portfolio optimization problems involving numerous assets and state-dependent dynamics has long been challenged by the curse of dimensionality. Traditional dynamic programming and PDE-based methods, while rigorous, typically become computationally intractable beyond a few state variables ($\sim$3-6 limit in prior studies). To overcome this critical barrier, we introduce the Pontryagin-Guided Direct Policy Optimization (PG-DPO) framework. PG-DPO leverages Pontryagin’s Maximum Principle (PMP) and backpropagation-through-time (BPTT) to directly inform neural network-based policy learning. A key contribution is our highly efficient Projected PG-DPO (P-PGDPO) variant. This approach uniquely utilizes BPTT to obtain rapidly stabilizing estimates of the Pontryagin costates and their crucial derivatives with respect to the state variables. These estimates are then analytically projected onto the manifold of optimal controls dictated by PMP’s first-order conditions, significantly reducing training overhead and enhancing accuracy. This enables a breakthrough in scalability: numerical experiments demonstrate that P-PGDPO successfully tackles problems with dimensions previously considered far out of reach (up to 50 assets and 10 state variables). Critically, the framework accurately captures complex intertemporal hedging demands, a feat often elusive for other methods in high-dimensional settings. P-PGDPO delivers near-optimal policies, offering a practical and powerful alternative for a broad class of high-dimensional continuous-time control problems.
@@ 1 Introduction
High-dimensional optimal control problems often fall prey to the “curse of dimensionality” (Bellman, , 1966), whereby the computational cost of classical dynamic programming (DP) approaches grows exponentially with the number of state variables. In a continuous-time finance setting, this difficulty becomes starkly evident in multi-asset Merton problems (Merton, , 1973), where an investor dynamically allocates wealth among numerous risky assets whose drift and volatility may depend on exogenous state variables such as macroeconomic indicators or stochastic volatility factors. While DP or Hamilton–Jacobi–Bellman (HJB) PDE solvers provide a rigorous framework for low-dimensional cases (Fleming and Soner, , 2006; Bertsekas and Tsitsiklis, , 1995), they become intractable as the number of assets or state dimensions increases. This difficulty is pervasive, impacting not only continuous-time settings (which are often addressed via discretization for numerical solution) but also analogous discrete-time models where similar state-space explosion issues arise. Consequently, prior influential studies in these domains are typically limited to analyzing problems involving up to approximately 3-5 state variables (Campbell and Viceira, , 1999; Lynch, , 2001; Balduzzi and Lynch, , 1999) or specific frameworks handling around 6 state variables (Garlappi and Skoulakis, , 2010), often necessitating simulation-based approaches for tractability (Brandt et al., , 2005).
@@
A promising alternative is to employ Pontryagin’s Maximum Principle (PMP), an indirect method that bypasses the explicit computation of value functions (Pontryagin, , 2014; Yong and Zhou, , 2012). Rather than storing or approximating a high-dimensional value grid, the PMP characterizes the optimal control in terms of forward–backward differential equations for the state and its costate (adjoint). Notably, adding one more state dimension does not produce an exponential explosion in computational requirements, in contrast to DP (Bellman, , 1966; Campbell et al., , 2003). This makes Pontryagin-based methods appealing for large-scale portfolio optimization problems involving many assets and exogenous state processes (Karatzas et al., , 1987).
@@
Recent advances in deep learning further enhance the scalability of methods for solving high-dimensional stochastic control problems (e.g., Han et al., , 2018; Raissi et al., , 2019, for general PDE-based approaches). Concurrently, deep learning techniques are being applied directly to financial dynamic programming and stochastic control. Some approaches tackle infinite horizon problems via the HJB equation (Duarte et al., , 2024), or solve the HJB equation for utility maximisation under constraints using formulations based on second-order backward stochastic differential equations (2BSDEs) (Davey and Zheng, , 2022). Others aim to directly solve associated nonlinear PDEs using methods like deep branching algorithms (Nguwi et al., , 2024). While these BSDE-based methods and direct PDE solvers are powerful, applications such as the Merton HJB have revealed limitations in accuracy or efficiency for certain variants (Nguwi et al., , 2024). Moreover, some works effectively leverage the Stochastic Maximum Principle (SMP) and adjoint BSDEs, often through duality methods, to handle even non-Markovian problems in finance (Davey and Zheng, , 2022). Alongside these PDE-centric or BSDE-centric deep learning approaches, another line of research, more akin to reinforcement learning, focuses on directly parameterizing the control policy with neural networks and optimizing the expected objective function through simulation and gradient ascent (e.g., Zhang and Zhou, , 2019; Kopeliovich and Pokojovy, , 2024, for portfolio problems with transaction costs or under Heston dynamics). Our PMP-based PG-DPO framework builds upon similar ideas of direct policy optimization but is distinctively guided by the continuous-time optimality conditions of PMP. It utilizes BPTT-derived costate estimates for finite-horizon problems, aiming for enhanced scalability and a novel way to enforce optimality conditions for problems with sufficient regularity.
One notable strategy related to the use of adjoint methods in optimal control, and central to our work, is to directly parameterize the policy (i.e., consumption and investment controls) with neural networks, then apply backpropagation-through-time (BPTT) to enforce Pontryagin’s optimality conditions. This strategy, which we term the baseline “Pontryagin-Guided Direct Policy Optimization (PG-DPO),” circumvents the need to approximate value functions or PDE solutions. Instead, it uses simulation-based gradients that automatically incorporate the adjoint processes, drastically reducing computational overhead in high-dimensional settings (Kushner and Yin, , 2003; Borkar and Borkar, , 2008).
@@
Furthermore, we introduce the “Projected PG-DPO” (P-PGDPO) variant, which significantly enhances efficiency and accuracy. This variant exploits the observation that BPTT-derived costate estimates often stabilize much faster than the full policy network. Once the costates are sufficiently stable (after a short warm-up phase), one can bypass the potentially suboptimal neural policy outputs. Instead, these costate estimates are directly substituted into the PMP first-order conditions (FOCs), effectively projecting the learned costate information onto the manifold of controls satisfying PMP’s necessary conditions. This recovers a near-optimal control analytically at every time-state point with minimal additional training, significantly reducing computational burden. While other deep learning methods such as Deep BSDE are also applied to high-dimensional problems, their ability to accurately capture complex, state-dependent optimal policy structures, particularly the crucial intertemporal hedging components, has been limited in challenging financial applications. Our numerical experiments (Section 4) demonstrate that P-PGDPO not only scales to problems with dozens of assets and state variables (e.g., n=50, k=10 in our experiments) but, critically, it is capable of precisely disentangling and reproducing these vital hedging demands—a feat not typically achieved by existing approaches in such high-dimensional settings. To the best of our knowledge, this explicit and accurate recovery of complex hedging terms in a high-dimensional, PMP-guided direct policy optimization framework is a novel contribution. This superior performance in both scalability and fidelity to the true optimal policy, especially its nuanced components, highlights the transformative potential of leveraging the PMP structure via BPTT-derived costate estimates for tackling a broad class of complex, high-dimensional control problems where amenability to costate estimation under suitable regularity conditions exists.
@@
  • We introduce Pontryagin-Guided Direct Policy Optimization (PG-DPO), and its highly efficient “Projected PG-DPO” (P-PGDPO) variant that uses PMP structure (specifically FOCs) and BPTT-derived costates to solve complex continuous-time control problems.
  • We demonstrate through numerical experiments that P-PGDPO not only breaks the dimensional barrier in portfolio optimization (up to n=50 assets, k=10 factors), but critically, accurately captures complex intertemporal hedging demands in these high-dimensional settings, a capability not typically demonstrated by alternative deep learning approaches.
In the remainder of this paper, we first present the multi-asset Merton problem with exogenous states in Section 2. Section 3 introduces our Pontryagin-Guided Direct Policy Optimization (PG-DPO) algorithm, its Projected PG-DPO (P-PGDPO) extension, and provides a theoretical justification for the effectiveness of the P-PGDPO approach under specified regularity conditions. Numerical experiments presented in Section 4 confirm the framework’s scalability and near-optimal performance within this testbed, illustrating its potential for high-dimensional applications. Finally, Section 5 concludes and discusses potential future directions.
@@ 2 Multi-Asset Continuous-Time Portfolio Problem with State Variables
In this section, we present a multi-asset extension of Merton’s continuous-time portfolio problem (Merton, , 1973; Kim and Omberg, , 1996; Liu, , 2007), allowing the investor to allocate wealth across multiple (potentially correlated) risky assets and a risk-free asset, while consuming continuously. Each risky asset features its own drift and volatility, leading to an $n$-dimensional portfolio weight vector $\bm{\pi}_{t}$. Furthermore, an exogenous state process $\mathbf{Y}_{t}\in\mathbb{R}^{k}$ influences market parameters such as returns and interest rates, adding additional complexity to the investor’s decision.
Although the wealth $X_{t}$ remains scalar, the investor’s optimal policy now depends on both $\bm{\pi}_{t}\in\mathbb{R}^{n}$ and a high-dimensional state $\mathbf{Y}_{t}$. We handle this through Pontryagin’s maximum principle (Pontryagin, , 2014; Yong and Zhou, , 2012), introducing a costate for $X_{t}$ and another for $\mathbf{Y}_{t}$. Numerically, one can represent $\{\bm{\pi}_{t},C_{t}\}$ via neural networks, thereby enabling policy gradient or dynamic programming algorithms even in multi-asset, state-dependent environments.
@@ 2.1 Problem Formulation with State Variables and Neural Network Controls
where $\bm{\mu}_{Y}(t,\mathbf{Y}_{t})$ is the $k\times 1$ drift vector, $\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})$ is a $k\times k$ diagonal matrix of diffusion coefficients (i.e., $\bm{\sigma}_{Y}(t,\mathbf{Y}_{t})=\operatorname{diag}(\sigma_{Y,1}(t,\mathbf{Y}_{t}),\dots,\sigma_{Y,k}(t,\mathbf{Y}_{t}))$ where $\sigma_{Y,i}$ are the diffusion functions from the component-wise definition), and $\mathbf{W}_{t}^{Y}$ is a $k$-dimensional Brownian motion whose components may be correlated.
@@
The investor’s objective is to choose an admissible control pair $(\bm{\pi},C)$ that maximises the expected discounted utility
[…]
subject to the wealth dynamics (1) and the admissibility constraints stated later in Assumption 2. Here $\delta>0$ is the subjective discount rate and $\kappa\geq 0$ weights the utility of terminal wealth (a bequest motive).
Throughout we work with the Hyperbolic Absolute Risk Aversion (HARA) family of utilities,
[…]
whose domain is the half-line $\mathcal{D}_{U}:=\{x>x_{\min}\}$ with $x_{\min}:=-a/b$. HARA nests the familiar special cases CRRA $(a=0,b=1)$ and CARA $(\gamma=1,b=1)$. On its domain $U$ is $C^{2}$, strictly increasing and concave, and it satisfies the Inada limits $U^{\prime}(x)\to\infty$ as $x\downarrow x_{\min}$ and $U^{\prime}(x)\to 0$ as $x\to\infty$ (see Pennacchi, , 2008, §3.2).
@@
    Interior bounds. There exist constants $\Pi_{\max}>0$ and $\xi\in(0,1)$ such that
@@
    Domain-admissible wealth. With $x_{\min}:=-a/b$ (see the HARA utility defined above) we require
@@
Assumption 2(iii) gives the strict inequalities $0<C_{t}<\xi X_{t}$, while Assumption 2(iv) guarantees the wealth process stays inside the utility’s domain, $X_{t}>x_{\min}\;(=-a/b)$, almost surely. Since $U$ is $C^{2}$, strictly concave, and satisfies the Inada limits on $(x_{\min},\infty)$, and because $\delta>0$, we obtain
$\mathbb{E}\!\left[e^{-\delta t}\,U(C_{t})\right]\;\leq\;e^{-\delta t}\,U\!\bigl{(}\xi\,\mathbb{E}[X_{t}]\bigr{)}\;<\;\infty,\qquad\mathbb{E}\!\bigl{[}U(X_{T})\bigr{]}<\infty,$
so the objective functional (2) is well defined for every admissible control pair $(\bm{\pi},C)$.
@@
    Uniform non-degeneracy of risky-asset volatility. Let $a(t,\mathbf{y}):=\bm{\sigma}(t,\mathbf{y})\,\Psi\,\bm{\sigma}(t,\mathbf{y})^{\top}$. There exists $\lambda>0$ such that
@@
    Utility specification. We use the HARA utility $U$ in (3); on its domain $x>x_{\min}$ it is $C^{2}$, strictly increasing ($U^{\prime}>0$), strictly concave ($U^{\prime\prime}<0$) and satisfies the Inada limits $U^{\prime}(x)\to\infty$ as $x\downarrow x_{\min}$ and $U^{\prime}(x)\to 0$ as $x\to\infty$.
@@
    Discount and bequest parameters. The subjective discount rate satisfies $\delta>0$, and the terminal-wealth weight satisfies $\kappa\geq 0$.
Part (a) ensures the diffusion is non-degenerate, which guarantees parabolic regularity of the associated Hamilton–Jacobi–Bellman (HJB) equation. Part (b) specialises preferences to the HARA family; it reduces to CRRA when $a=0$ (irrespective of $b$), and to the familiar log-utility when $a=0$ and $\gamma=1$. Together with the coefficient regularity of Assumption 1, these conditions yield existence, uniqueness, and sufficient smoothness of the value function (e.g. $V\in C^{1,2,2}$); see (Oksendal, , 2013, Chs. 6–7).